OpenStudy (anonymous):

How do you solve a problem that looks like this: 11x^4 + 8x^3 - 2x^2 + 39 ?

OpenStudy (anonymous):

ok let me start from the beggining

OpenStudy (anonymous):

\[11x ^{4}+8x ^{3}-2x ^{2}39\]

OpenStudy (anonymous):

oh, damn, i gota go eat, my mom is calling.... sorry

OpenStudy (anonymous):

ok...just answer it later I guess

OpenStudy (anonymous):

This problem has no solution

OpenStudy (anonymous):

it's not that problem I'm trying to solve, just asking how to solve a problem like it. like how I isolate the variables

OpenStudy (anonymous):

There is no specific rule

OpenStudy (anonymous):

you want to simplify?

OpenStudy (anonymous):

There is one rule of finding the discriminant. But I am not sure if they are within the scope of your syllabus

OpenStudy (anonymous):

sstarica, what I'm trying to do is solve for x.

OpenStudy (amistre64):

11x^4 + 8x^3 - 2x^2 + 39; does it equal 0? otherwise it can be any number.. Lets assume that to solve for the "roots" of this equation we are making it equal to zero.. factor out an x^2 from the first 3 terms: (x^2)(11x^2 + 8x - 2) + 39 = 0 factor the quadratic: 11x^2 + 8x - 2 sqrt(64 - (4)(-2)(11)) = sqrt(64+88) = (-8/22) +- (sqrt(152)/22) x= (-4/11) + sqrt(152)/22 or (-4/11) - sqrt(152)/22 possibly :)

OpenStudy (amistre64):

or at least thats as far as I got... gotta take into account the +39 and the x^2 stuff

OpenStudy (amistre64):

could try grouping and see if it works it out....